# Weighted least-squares approximation with determinantal point processes and generalized volume sampling

### 21 Dec 2023 v1

@@ 5.1 Independent repetitions of DPP distribution $\gamma_{m}$
Let $\mathbf{z}\sim\gamma_{m}$ and $\mathbf{y}\sim\nu_{m}^{\otimes m}$. Then
$\mathbb{E}(F(\mathbf{A}(\mathbf{z})))\leq\mathbb{E}(F(\mathbf{A}(\mathbf{y})))$ (13)
where $F$ is a real-valued positive and monotonically decreasing function in the Loewner order.
If 5.3 is verified, we obtain the following result.
@@
Let $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{r})\sim\Gamma_{m}^{\otimes r}$ where $\Gamma_{m}$ is a distribution over $\mathcal{X}^{m}$ satisfying
$\mathbb{E}_{\mathbf{z}\sim\Gamma_{m}}(F(\mathbf{A}(\mathbf{z})))\leq\mathbb{E}_{\mathbf{y}\sim\nu_{m}^{\otimes m}}(F(\mathbf{A}(\mathbf{y})))$
for $F$ a real-valued positive and monotonically decreasing function in the Loewner order. Then
@@
We have $\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{x}))<1-\delta)=\mathbb{E}(F(\mathbf{G}^{w_{m}}(\mathbf{x})))$ with $F(\mathbf{B})=\mathbf{1}_{\|\mathbf{B}^{-1}\|_{2}>(1-\delta)^{-1}}$. If $\mathbf{B}\preceq\mathbf{C}$, then $\mathbf{C}^{-1}\preceq\mathbf{B}^{-1}$, $\|\mathbf{C}^{-1}\|_{2}\leq\|\mathbf{B}^{-1}\|_{2}$ and therefore, $F(\mathbf{C})\geq F(\mathbf{B})$. That means $F$ is decreasing in the Loewner order. Letting $\mathbf{G}^{w_{m}}=\frac{1}{r}\sum_{i=1}^{m}\mathbf{A}(\mathbf{x}_{i}):=\mathbf{H}(\mathbf{x}_{1},\ldots,x_{r-1})+\mathbf{A}(\mathbf{x}_{r})/r$, and noting that $\mathbf{A}\mapsto F(\mathbf{H}+\mathbf{A})$ is decreasing in the Loewner order, it holds
$\mathbb{E}(F(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i})))$ $=\mathbb{E}(\mathbb{E}(F(\mathbf{H}(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1})+\mathbf{A}(\mathbf{x}_{r})/r)|\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1}))$
$\leq\mathbb{E}(F(\mathbf{H}(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1})+\mathbf{A}(\mathbf{y}_{r})/r)),$
and we proceed recursively to obtain
$\mathbb{E}(F(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i})))$ $\leq\mathbb{E}(F(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{y}_{i}))),$
where the $\mathbf{y}_{i}$ are i.i.d. samples from $\nu_{m}^{\otimes m}.$ This yields $\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{x}))<1-\delta)\leq\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{y}))<1-\delta)$, with $\mathbf{y}\sim\nu^{\otimes n}$, $n=rm$, and we conclude using Lemma 3.1. ∎
@@ 5.2 Independent repetitions of volume sampling $\gamma^{\nu}_{\bar{n}}$
A better understanding of the distribution of matrices $\mathbf{G}^{w}(\mathbf{x}_{k})$ allows to improve the above results. As for the case of determinantal point processes, we conjecture that $\mathbb{E}(F(\mathbf{G}^{w}(\mathbf{x}_{k})))\leq\mathbb{E}_{\mathbf{y}\sim\nu^{\otimes\bar{n}}}(F(\mathbf{G}^{w}(\mathbf{y})))$ for $F$ a real-valued positive and monotonically decreasing function in the Loewner order. Under this conjecture, following the proof of Proposition 5.4, we would obtain the following concentration result, similar to the case of $n=m\bar{n}$ i.i.d. sampling from $\nu=w\mu$.
@@
$\mathbb{E}_{\mathbf{z}\sim\Gamma^{\nu}_{\bar{n}}}(F(\mathbf{G}^{w}(\mathbf{z})))\leq\mathbb{E}_{\mathbf{y}\sim\nu^{\otimes\bar{n}}}(F(\mathbf{G}^{w}(\mathbf{y})))$
for $F$ a real-valued positive and monotonically decreasing function in the Loewner order. Then it holds
@@
We can obtain a similar result by assuming that the 5.3 on the projection DPP holds, and further assuming that $w\leq\xi_{m}w_{m}$ for some constant $\xi_{m}$ (possibly depending on $m$).
@@
Let $\bar{n}\geq m$, $r\in\mathbb{N}$ and $n=\bar{n}m$. Let $\mathbf{y}\sim\nu^{\otimes(n-mr)}$ with $\nu=w\mu$ such that $w\leq\xi_{m}w_{m}$. Let $\mathbf{z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{r})\sim\Gamma_{m}^{\otimes r}$ where $\Gamma_{m}$ is a distribution over $\mathcal{X}^{m}$ satisfying
$\mathbb{E}_{\mathbf{z}\sim\Gamma_{m}}(F(\mathbf{A}(\mathbf{z})))\leq\mathbb{E}_{\mathbf{y}\sim\nu_{m}^{\otimes m}}(F(\mathbf{A}(\mathbf{y})))$
for $F$ a real-valued positive and monotonically decreasing function in the Loewner order. Letting $\mathbf{x}=(\mathbf{y},\mathbf{z})$, it holds
@@
$\mathbf{G}^{w}(\mathbf{x})=\frac{mr}{n}\mathbf{G}^{w}(\mathbf{z})+\frac{n-mr}{n}\mathbf{G}^{w}(\mathbf{y})\succeq\frac{mr}{n}\mathbf{G}^{w}(\mathbf{z})\succeq\frac{mr}{n}\xi_{m}^{-1}\mathbf{G}^{w_{m}}(\mathbf{z}),$
and therefore, using Proposition 5.4,

### 30 Jan 2024 v2

@@ 5.1 Independent repetitions of DPP distribution $\gamma_{m}$
The distribution $\gamma_{m}$ satisfies
$\mathbb{P}_{\mathbf{z}\sim\gamma_{m}}(F(\mathbf{A}(\mathbf{z}))>t)\leq\mathbb{P}_{\mathbf{y}\sim\nu_{m}^{\otimes m}}(F(\mathbf{A}(\mathbf{y}))>t)$ (13)
for $t>0$ and $F$ a real-valued positive, convex and decreasing function in the Loewner order.
If the distribution $\gamma_{m}$ satisfies the property of 5.3, we obtain the following result.
@@
Let $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{r})\sim\gamma_{m}^{\otimes r}$ with $\gamma_{m}$ a distribution over $\mathcal{X}^{m}$ satisfying (13). Then
@@
We have $\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{x}))<1-\delta)=\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{x}))^{-1}>t)$ with $t:=(1-\delta)^{-1}$. The function $F:=\mathbf{B}\mapsto\lambda_{min}(\mathbf{B})^{-1}$ is a positive convex and monotonically decreasing function in the Loewner order. For any fixed symmetric positive semi-definite matrix $\mathbf{H}$, $\mathbf{A}\mapsto\lambda_{min}(\mathbf{H}+\mathbf{A}/r)^{-1}$ is also a positive convex and monotonically decreasing function in the Loewner order. Letting $\mathbf{G}^{w_{m}}(\mathbf{x})=\frac{1}{r}\sum_{i=1}^{m}\mathbf{A}(\mathbf{x}_{i}):=\mathbf{H}(\mathbf{x}_{1},\ldots,x_{r-1})+\mathbf{A}(\mathbf{x}_{r})/r$, we then have
$\mathbb{P}(F(\mathbf{G}^{w_{m}}(\mathbf{x}))>t)$ $=\mathbb{E}(\mathbf{1}_{F(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i}))>t})$
$=\mathbb{E}(\mathbb{E}(\mathbf{1}_{F(\mathbf{H}(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1})+\mathbf{A}(\mathbf{x}_{r})/r)>t}|\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1}))$
$\leq\mathbb{E}(\mathbb{E}(\mathbf{1}_{F(\mathbf{H}(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1})+\mathbf{A}(\mathbf{y}_{r})/r)>t}|\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1}))$
$=\mathbb{E}(\mathbf{1}_{F(\mathbf{H}(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-1})+\mathbf{A}(\mathbf{y}_{r})/r)>t}),$
where $\mathbf{y}_{r}\sim\nu_{m}^{\otimes m}$. Letting $\mathbf{y}_{1},\ldots,\mathbf{y}_{r}$ be i.i.d. samples from $\nu_{m}^{\otimes m}$, and successively conditioning on $(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-2},\mathbf{y}_{r})$, $(\mathbf{x}_{1},\ldots,\mathbf{x}_{r-3},\mathbf{y}_{r-1},\mathbf{y}_{r})$, …, $(\mathbf{y}_{2}\ldots,\mathbf{y}_{r})$, we obtain
$\mathbb{P}(F(\mathbf{G}^{w_{m}}(\mathbf{x}))>t)$ $\leq\mathbb{E}(\mathbf{1}_{F(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{y}_{i}))>t})=\mathbb{P}(F(\mathbf{G}^{w_{m}}(\mathbf{y}))>t),$
where $\mathbf{y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{r})\sim\nu_{m}^{\otimes n}$, $n=rm$. Then it holds $\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{x}))<1-\delta)\leq\mathbb{P}(\lambda_{min}(\mathbf{G}^{w_{m}}(\mathbf{y}))<1-\delta)$, and we conclude using Lemma 3.1. ∎
@@
Remark 5.5.
The proof of Proposition 5.4 exploits the assumption (13) to prove that
$\mathbb{P}_{\mathbf{z}\sim\gamma_{m}}(\lambda_{min}(\mathbf{H}+\mathbf{A}(\mathbf{z})/r)^{-1}>t)\leq\mathbb{P}_{\mathbf{y}\sim\nu_{m}^{\otimes m}}(\lambda_{min}(\mathbf{H}+\mathbf{A}(\mathbf{y})/r)^{-1}>t)$ (14)
for any fixed p.s.d. matrix $\mathbf{H}$. The assumption (14) on $\gamma_{m}$ would be sufficient to obtain the result of Proposition 5.4.
Remark 5.6.
In order to obtain the result of Proposition 5.4, an alternative assumption on $\gamma_{m}$ would be that
$\mathbb{E}_{\mathbf{z}\sim\gamma_{m}}(G(e^{s\mathbf{A}(\mathbf{z})}))\leq\mathbb{E}_{\mathbf{y}\sim\nu_{m}^{\otimes m}}(G(e^{s\mathbf{A}(\mathbf{y})}))$ (15)
for any $s<0$ and $G$ a real-valued positive, concave and monotonically increasing function in the Loewner order. Under this assumption, we have to follow the proof of matrix Chernoff. The first steps of the proof of matrix Chernoff inequality (Theorem A.1) yield
$\mathbb{P}(\lambda_{min}(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i}))<t)\leq\inf_{\theta<0}e^{-\theta t}\mathbb{E}(\mathrm{tr}\exp(\sum_{i=1}^{r}\theta\mathbf{A}(\mathbf{x}_{i})/r)).$
Letting $\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i}):=\mathbf{H}+\mathbf{A}(\mathbf{x}_{r})/r$, we have
$\mathrm{tr}\exp(\sum_{i=1}^{r}\theta\mathbf{A}(\mathbf{x}_{i})/r)=G(e^{\mathbf{A}(\mathbf{x}_{r})\theta/r})$
with $G:=\mathbf{X}\mapsto\mathrm{tr}\exp(\mathbf{H}+\log(\mathbf{X}))$ a concave and increasing function in the Loewner order. The assumption then implies that
$\mathbb{E}(\mathrm{tr}\exp(\sum_{i=1}^{r}\theta\mathbf{A}(\mathbf{x}_{i})/r))\leq\mathbb{E}(\mathrm{tr}\exp(\mathbf{H}+\mathbf{A}(\mathbf{y}_{r})/r))$
with $\mathbf{y}_{r}\sim\nu_{m}^{\otimes m}$. Then by successive conditioning (as in the proof of Proposition 5.4), we obtain
$\mathbb{P}(\lambda_{min}(\frac{1}{r}\sum_{i=1}^{r}\mathbf{A}(\mathbf{x}_{i}))<t)\leq\inf_{\theta<0}e^{-\theta t}\mathbb{E}(\mathrm{tr}\exp(\sum_{i=1}^{r}\theta\mathbf{A}(\mathbf{y}_{i})/r)),$
where the $\mathbf{y}_{i}$ are i.i.d. samples from $\nu_{m}^{\otimes m}$, and we proceed with the classical proof of matrix Chernoff inequality for sums of i.i.d. matrices.
@@ 5.2 Independent repetitions of volume sampling $\gamma^{\nu}_{\bar{n}}$
A better understanding of the distribution of matrices $\mathbf{G}^{w}(\mathbf{x}_{k})$ allows to improve the above results. As for the case of determinantal point processes, we conjecture that
$\mathbb{P}(F(\mathbf{G}^{w}(\mathbf{x}_{k}))>t)\leq\mathbb{P}_{\mathbf{y}\sim\nu^{\otimes\bar{n}}}(F(\mathbf{G}^{w}(\mathbf{y}))>t)$ (16)
for $t>0$ and $F$ a real-valued positive, convex and monotonically decreasing function in the Loewner order. Under this conjecture, following the proof of Proposition 5.4, we would obtain the following concentration result, similar to the case of $n=m\bar{n}$ i.i.d. sampling from $\nu=w\mu$.
@@
Remark 5.11.
The assumption (16) in Proposition 5.10 could be replaced by a weaker condition of the form (14), or an alternative condition of the form (15).
We can avoid any conjecture on volume sampling and still obtain a result similar to Proposition 5.10 by assuming that the DPP distribution $\gamma_{m}$ satisfies the conjecture (13) (or one of the two conjectures (14) or (15)), and further assuming that $w\leq\xi_{m}w_{m}$ for some constant $\xi_{m}$ (possibly depending on $m$).
@@
Let $\bar{n}\geq m$, $r\in\mathbb{N}$ and $n=\bar{n}m$. Let $\mathbf{y}\sim\nu^{\otimes(n-mr)}$ with $\nu=w\mu$ such that $w\leq\xi_{m}w_{m}$. Let $\mathbf{z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{r})\sim\gamma_{m}^{\otimes r}$ where $\gamma_{m}$ is a distribution over $\mathcal{X}^{m}$ satisfying either (13), (14) or (15). Letting $\mathbf{x}=(\mathbf{y},\mathbf{z})$, it holds
@@
$\mathbf{G}^{w}(\mathbf{x})=\frac{mr}{n}\mathbf{G}^{w}(\mathbf{z})+\frac{n-mr}{n}\mathbf{G}^{w}(\mathbf{y})\succeq\frac{mr}{n}\mathbf{G}^{w}(\mathbf{z})\succeq\frac{mr}{n}\xi_{m}^{-1}\mathbf{G}^{w_{m}}(\mathbf{z}).$
Therefore, using Proposition 5.4 with assumption (13) or the alternative assumptions (14) or (15) (see Remarks 5.5 and 5.6), it holds
@@ Appendix A Some known results on random matrices
Proof.
We provide a sketch of the proof of Tropp [tropp2015introduction] for the bound on the minimal eigenvalue. Let $\mathbf{B}=\sum_{i=1}^{n}\mathbf{A}_{i}$. For any $\theta<0$, it holds
$\mathbb{P}(\lambda_{min}(\mathbf{B})\leq t)=\mathbb{P}(e^{\theta\lambda_{min}(\mathbf{B})}\geq e^{\theta t})=\mathbb{P}(e^{\lambda_{min}(\theta\mathbf{B})}\geq e^{\theta t})\leq e^{-\theta t}\mathbb{E}(e^{\lambda_{min}(\theta\mathbf{B})}),$
where the last inequality is given by Markov inequality. Then using $e^{\lambda_{min}(\theta\mathbf{B})}=\lambda_{min}(e^{\theta\mathbf{B}})\leq\mathrm{tr}(e^{\theta\mathbf{B}})$, we obtain
$\mathbb{P}(\lambda_{min}(\mathbf{B})\leq t)\leq\inf_{\theta<0}e^{-\theta t}\mathbb{E}(\mathrm{tr}e^{\theta\mathbf{B}}).$
For positive-definite matrix $\mathbf{H}$, the map $\mathbf{A}\mapsto\mathrm{tr}e^{\mathbf{H}+\log(\mathbf{A})}$ is concave on the positive cone of positive definite matrices. Letting $\mathbf{X}_{i}:=e^{\theta\mathbf{A}_{i}}$, a sequential application of Jensen’s inequality gives
$\mathbb{E}(\mathrm{tr}e^{\theta\mathbf{B}})=\mathbb{E}(\mathrm{tr}\exp(\sum_{i=1}^{n}\log(\mathbf{X}_{i})))\leq\mathrm{tr}\exp(\sum_{i=1}^{n}\log(\mathbb{E}(\mathbf{X}_{i})))=\mathrm{tr}\exp(\sum_{i=1}^{n}\log(\mathbb{E}(e^{\theta\mathbf{A}_{i}}))).$
Also it holds $\log\mathbb{E}(e^{\theta\mathbf{A}_{i}})\preceq g(\theta)\mathbb{E}(\mathbf{A}_{i})$ where $g(\theta)=L^{-1}(e^{\theta L}-1),$ so that
$\mathrm{tr}\exp(\sum_{i=1}^{n}\log(\mathbb{E}(e^{\theta\mathbf{A}_{i}})))\leq\mathrm{tr}\exp(g(\theta)\mathbb{E}(\sum_{i=1}^{n}\mathbf{A}_{i}))\leq ne^{g(\theta)\mu_{min}}.$
Therefore,
$\mathbb{P}(\lambda_{min}(\mathbf{B})\leq t)\leq\inf_{\theta<0}ne^{-\theta t}e^{g(\theta)\mu_{min}}$
Taking $t=(1-\delta)\mu_{min}$, the infimum is attained at $\theta=L^{-1}\log(1-\delta)$, which gives the desired result. ∎